Q6. [20 marks] The estimated annual cash flows of an investment project along with associated probabilities are given below.ID is even
Answer #1

Answer:

Given Data

i = 12%

Expected PW = 0.45*PW1+ 0.25 * PW2 + 0.3 *PW3

PW_1= -10000+4200/1.12+3250/1.12^{2}+--3250/1.12^{6}+6000/1.12^{7} = -10000+16923.96

= 6923.96

PW_2 = -20000+3500/1.12+3500/1.12^{2}+--3500/1.12^{6}+5000/1.12^{7} = -20000+ 16651.67

= -3348.33

PW_3=-12000+3100/1.12+3100/1.12^2+--3100/1.12^6+3100/1.12^7 = -12000+14147.623

= -2147.623

Expected PW =   0.45*PW1+ 0.25 * PW2 + 0.3 *PW3

= 0.45(6923.96) + 0.25 (-3348.33)+0.3( -2147.623)

= 3115.782 - 837.0825 -536.90575

= 1741.79375